OpenStudy (anonymous):

3 + sqrt z - 1 = sqrt z + 14

OpenStudy (anonymous):

as it is posted, the question is a contradiction. post it clearly. either draw your question or use the equation button below.

OpenStudy (anonymous):

\[3+\sqrt{z}-1=\sqrt{z}+14\] \[2+\sqrt{z}-\sqrt{z}=14\] \[2=14 \] that is not possible hence no solution